A fair coin is tesse, and ה fair sx sided die 1s rolled. Suppose the outcomes for the c ein and die are Tie endent. Te pmbibilitv of geting a need an roling a six is A. 112 B 23 C 13 2. Suppose we roll a red die and a gren dic. LetAbe the ecl thal he nanber of spots showing en lbe red die the1 the number of spots shwing on the green die is more than thee. Thesents A and B are: three or less and B be the cveni A.disjomtcamplements. C.indeperdentD.recipocals EventAccurs with probabiliy 02. vent B oceurs with probabiliy D8. If A and Bnre disiont utually exclusive, then: A, P(A and B)-0.16. B.PiAcr B)-l.0 C.PiA anc B)-1.ต. D. PA orB)-0.16. 1. Suppose thbt A adB ar two independent events with PA)-0.3 P-0.3 PIAor B is: 5 Yo are making a DIY cookie. There are kinds of flours. 4 kinis af flavour, and 5 kind of n卩ping How many kinds ofonkies can yau nake? 6. 2 students are running f the coemmittee election, ard you need a arrange their order for election speech 201 students are running frelecti and only of them wwal he a:littelt the cmmittee. 8. there are 2l studeats in he committee, and we are selecting3 students to he President, VP and Finance. Page
Answer #1

Solution:-

1) (A) The probability of getting a head and rolling a six is 1/12.

The probability of getting a head = 1/2

The probability of rolling a six = 1/6

The probability of getting a head and rolling a six = (1/2)*(1/6)

The probability of getting a head and rolling a six is 1/12.

2) (C) Independent.

3) (B) P(A or B) = 1

For Mutually exclusive events P(A and B) = 0

P(A or B) = P(A) + P(B)

P(A or B) = 1